A circuit containing Resistance, Inductance and capacitance of 20 ohm, 10mH, 300µF are connected in series respectively ,across a supply voltage of v=100sin100Πt. Find out the expression of current


1
Expert's answer
2021-05-18T08:37:04-0400

The expression for current is


"i=\frac vz,\\\space\\
z=\sqrt{(R^2+(X_L-X_C)^2)},\\\space\\
X_L=\omega L,X_C=1/(\omega C).\\\space\\
i=\frac{100\sin(100\pi t)}{\sqrt{20^2+(0.01\cdot100\pi-1/(0.0003\cdot100π))^2}}.\\\space\\
i(t)=4.68\sin(100πt)."